# Estimating the Heritability of Longitudinal Rate-of-Change: Genetic Insights into PSA Velocity in Prostate Cancer-Free Individuals

### v1

@@ 2.1 Estimating Heritability in Cross-Sectional Studies
The covariance matrix for $\hat{\boldsymbol{\theta}}$ is obtained as $-\boldsymbol{AI}_{\boldsymbol{\theta}}(\hat{\boldsymbol{\theta}})^{-1}$ after convergence and the variance of $\hat{h}^{2}$ can be derived by the Delta method (Lynch and Walsh, 1998; Lehmann and Casella, 2006). Of note, $\boldsymbol{V}$ can be expressed as a linear combination of multiple variance components, a necessary condition for the AI-REML algorithm (Johnson and Thompson, 1995).
@@ 2.2.1 Leveraging AI-REML Algorithm for Studies with Small-to-Moderate Sample Sizes
[…]
@@
$-\boldsymbol{AI}_{\boldsymbol{\xi}}(\boldsymbol{\hat{\xi}})\approx-\hat{\boldsymbol{K}}^{\top}\boldsymbol{AI}_{\boldsymbol{\theta}}(\hat{\boldsymbol{\theta}})\hat{\boldsymbol{K}},$ (10)
where $\boldsymbol{AI}(\hat{\boldsymbol{\theta}})$ is the average information matrix for $\boldsymbol{\theta}$ evaluated at $\hat{\boldsymbol{\theta}}$. The covariance matrix for $(\hat{\lambda}_{1},\hat{\lambda}_{2})^{\top}$ can be extracted from the inverse of $-\boldsymbol{AI}_{\boldsymbol{\xi}}(\boldsymbol{\hat{\xi}})$.
@@ 3 Simulation Studies
Table 1: Parameter settings used in Scenarios I–III, based on $1{,}000$ repetitions, where $N=2{,}000$ (or $15{,}260$), $J=6$ (or $10$), and $P=10{,}000$ (or $6{,}948{,}674$).
@@
[…]
@@
Mean is the average of estimates from 1,000 repetitions; Median is the median of estimates from 1,000 repetitions; SE is the estimated standard error derived from estimated Fisher’s information; Empirical SE is calculated across repetitions; MAD represents the scaled median absolute deviation so that it corresponds to the SE under Gaussian distribution and provides a robust measure of variability; A dash () indicates that SE was not available for the REHE method
@@
[…]
@@
Mean is the average of estimates from 1,000 repetitions; Median is the median of estimates from 1,000 repetitions; SE is the estimated standard error derived from estimated Fisher’s information; Empirical SE is calculated across repetitions; MAD represents the scaled median absolute deviation so that it corresponds to the SE under Gaussian distribution and provides a robust measure of variability; A dash () indicates that SE was not available for the REHE method
@@ 4 Application to PSA Analysis
A direct application of AI-REML for fitting Model (15) would require the inversion of a $454\times 81,$ 454 × 81 , 454 dimensional covariance matrix since our dataset has $15,260$ participants with an average of $5.34$ measurements per individual; this makes the estimation computationally infeasible. We randomly partition the population into five groups, each containing $3,052$ males, and then apply the AI-REML algorithm on each subsample. Variance components and the two derived heritability metrics are separately estimated for each partition. Table 4 presents the partition-specific estimates from the AI-REML algorithm (columns 2-6); note that in some cases, these estimates are at the boundary of the parameter space (this only occurs when estimating the variance components and heritability related to slope). To obtain combined estimates of the variance components ($\boldsymbol{\theta}$) and two heritability metrics ($\lambda_{1}$ and $\lambda_{2}$) and their standard errors, we performed truncation-adjusted meta-analysis. For each variance component, we used left-truncated adjustment since there is only a lower bound at zero and no upper bound. For the heritability metrics ($\lambda_{1}$ and $\lambda_{2}$), those of which is bounded by 0 and 1, we used a doubly truncated adjustment. The results indicate substantial heritability for both the intercept and slope, estimated at 0.31 (SE = 0.07) and 0.45 (SE = 0.18), respectively. As a sensitivity analysis, we performed the analysis by partitioning the population into seven groups and obtained similar results (Supplementary Material D). As seen in the simulation results, standard errors for estimating slope heritability are substantially larger than those for the intercept, demonstrating the difficulty in making inference on the genetic influences for velocity. In the PLCO Study, we had an average number of follow-up measurements across participants of $5.34$. With additional follow-up measurements, we would certainly decrease the standard error for $\lambda_{2}$. In the simulations, we showed that there was a substantial decrease in the standard errors for $\lambda_{2}$ when we increased the number of follow-up measurements per subject from 6 to 10 (Table 3). Moreover, the two estimated heritability metrics, combined from partition-specific estimates using the simple averaging method or fixed-effect meta-analysis, were comparable to those derived from the truncation-adjusted meta-analysis (Supplementary Material E). This consistency is expected, as our simulation studies indicated minimal bias in estimating $\lambda_{2}$ when this heritability metric is approximately $0.50$.
@@
Table 5 shows the results when we applied the REHE method to the PLCO Study data, we find that the estimated ratios of genetic contributions to the baseline PSA levels and the rate of change in PSA level are $0.26\ (\text{SE}=0.06)$ and $0.48\ (\text{SE}=0.37)$, respectively. A comparison of the AI-REML algorithm and REHE method shows similar estimates of heritability metrics and standard errors for heritability on the intercept ($\lambda_{1}$), but notable differences in the standard errors for velocity heritability ($\lambda_{2}$). This observation is consistent with findings from the simulation studies, where AI-REML, particularly when combined with truncation-adjusted meta-analysis, demonstrated superior performance over REHE, especially in estimating the velocity heritability. Overall, based on the AI-REML results for the longitudinal PSA data from the PLCO study, the heritability of baseline PSA was estimated at $0.32\ (\text{SE}=0.07)$ and the heritability of PSA velocity at $0.45\ (\text{SE}=0.18)$, suggesting a substantial genetic contribution to PSA velocity. This finding highlights the potential for genome-wide association studies of PSA velocity and further genetically adjusted PSA velocity to enhance the efficacy of PSA screening for the early detection of (aggressive) prostate cancer.

### v2

@@ 2.1 Estimating Heritability in Cross-Sectional Studies
The covariance matrix for $\hat{\boldsymbol{\theta}}$ is obtained as $\boldsymbol{AI}_{\boldsymbol{\theta}}(\hat{\boldsymbol{\theta}})^{-1}$ after convergence and the variance of $\hat{h}^{2}$ can be derived by the Delta method (Lynch and Walsh, 1998; Lehmann and Casella, 2006). Of note, $\boldsymbol{V}$ can be expressed as a linear combination of multiple variance components, a necessary condition for the AI-REML algorithm (Johnson and Thompson, 1995).
@@ 2.2.1 Leveraging AI-REML Algorithm for Studies with Small-to-Moderate Sample Sizes
We use the Delta method to obtain the covariance matrix $\text{Cov}((\hat{\lambda}_{1},\hat{\lambda}_{2})^{\top})$. To proceed, we define $\boldsymbol{\xi}=(\xi_{1},\xi_{2},\xi_{3},\xi_{4},\xi_{5})^{\top}:=(\lambda_{1},\lambda_{2},\sigma^{2}_{g}+\sigma^{2}_{b0},\sigma^{2}_{g^{*}}+\sigma^{2}_{b1},\sigma^{2}_{e})^{\top}$. In Supplementary Material B, we derive the Jacobian matrix
@@
$\boldsymbol{AI}_{\boldsymbol{\xi}}(\boldsymbol{\hat{\xi}})\approx-\hat{\boldsymbol{K}}^{\top}\boldsymbol{AI}_{\boldsymbol{\theta}}(\hat{\boldsymbol{\theta}})\hat{\boldsymbol{K}}$, (10)
[…]
@@ 3 Simulation Studies
Table 1: Parameter settings used in Scenarios I–III, based on 1,0001{,}000 repetitions, where N=2, [^0] N=2{,}000 (or 15,26015{,}260), $J=6$ (or $10$), and P=10, [^0] P=10{,}000 (or 6,948,6746{,}948{,}674).
@@
Mean, Median(SE,Emp SE,MAD)
$N=2,000$
Parameter Scenario True Method $P=10,000,J=6$ $P=10,000,J=10$ $P=6,948,674,J=6$ $P=6,948,674,J=10$
@@
Mean is the average of estimates from 1,000 repetitions; Median is the median of estimates from 1,000 repetitions; SE is the estimated standard error derived from estimated Fisher’s information; Empirical SE is calculated across repetitions; MAD represents the scaled median absolute deviation so that it corresponds to the SE under Gaussian distribution and provides a robust measure of variability; A dash (—) indicates that SE was not available for the REHE method
@@
Mean, Median(SE,Emp SE,MAD)
$N=2,000$
Parameter Scenario True Method $P=10,000,J=6$ $P=10,000,J=10$ $P=6,948,674,J=6$ $P=6,948,674,J=10$
@@
Mean is the average of estimates from 1,000 repetitions; Median is the median of estimates from 1,000 repetitions; SE is the estimated standard error derived from estimated Fisher’s information; Empirical SE is calculated across repetitions; MAD represents the scaled median absolute deviation so that it corresponds to the SE under Gaussian distribution and provides a robust measure of variability; A dash (—) indicates that SE was not available for the REHE method
@@ 4 Application to PSA Analysis
A direct application of AI-REML for fitting Model (15) would require the inversion of a $81,454\times 81,454$ dimensional covariance matrix since our dataset has $15,260$ participants with an average of $5.34$ measurements per individual; this makes the estimation computationally infeasible. We randomly partition the population into five groups, each containing $3,052$ males, and then apply the AI-REML algorithm on each subsample. Variance components and the two derived heritability metrics are separately estimated for each partition. Table 4 presents the partition-specific estimates from the AI-REML algorithm (columns 2-6); note that in some cases, these estimates are at the boundary of the parameter space (this only occurs when estimating the variance components and heritability related to slope). To obtain combined estimates of the variance components ($\boldsymbol{\theta}$) and two heritability metrics ($\lambda_{1}$ and $\lambda_{2}$) and their standard errors, we performed truncation-adjusted meta-analysis. For each variance component, we used left-truncated adjustment since there is only a lower bound at zero and no upper bound. For the heritability metrics ($\lambda_{1}$ and $\lambda_{2}$), those of which is bounded by 0 and 1, we used a doubly truncated adjustment. The results indicate substantial heritability for both the intercept and slope, estimated at 0.31 (SE = 0.07) and 0.45 (SE = 0.18), respectively. As a sensitivity analysis, we performed the analysis by partitioning the population into seven groups and obtained similar results (Supplementary Material D). As seen in the simulation results, standard errors for estimating slope heritability are substantially larger than those for the intercept, demonstrating the difficulty in making inference on the genetic influences for velocity. In the PLCO Study, we had an average number of follow-up measurements across participants of $5.34$. With additional follow-up measurements, we would certainly decrease the standard error for $\lambda_{2}$. In the simulations, we showed that there was a substantial decrease in the standard errors for $\lambda_{2}$ when we increased the number of follow-up measurements per subject from 6 to 10 (Table 3). Moreover, the two estimated heritability metrics, combined from partition-specific estimates using the simple averaging method or fixed-effect meta-analysis, were comparable to those derived from the truncation-adjusted meta-analysis (Supplementary Material E). This consistency is expected, as our simulation studies indicated minimal bias in estimating $\lambda_{2}$ when this heritability metric is approximately $0.50$.
@@
Table 5 shows the results when we applied the REHE method to the PLCO Study data, we find that the estimated ratios of genetic contributions to the baseline PSA levels and the rate of change in PSA level are 0.26(SE=0.06)0.26\ ($\text${SE}=0.06) and 0.48(SE=0.37)0.48\ ($\text${SE}=0.37), respectively. A comparison of the AI-REML algorithm and REHE method shows similar estimates of heritability metrics and standard errors for heritability on the intercept ($\lambda_{1}$), but notable differences in the standard errors for velocity heritability ($\lambda_{2}$). This observation is consistent with findings from the simulation studies, where AI-REML, particularly when combined with truncation-adjusted meta-analysis, demonstrated superior performance over REHE, especially in estimating the velocity heritability. Overall, based on the AI-REML results for the longitudinal PSA data from the PLCO study, the heritability of baseline PSA was estimated at 0.32(SE=0.07)0.32\ ($\text${SE}=0.07) and the heritability of PSA velocity at 0.45(SE=0.18)0.45\ ($\text${SE}=0.18), suggesting a substantial genetic contribution to PSA velocity. This finding highlights the potential for genome-wide association studies of PSA velocity and further genetically adjusted PSA velocity to enhance the efficacy of PSA screening for the early detection of (aggressive) prostate cancer.